Solve for t. Give an exact answer. ln(t)-ln(t-4)=1
Answer by jim_thompson5910(35256)

ln(t)-ln(t-4)=1

ln(t/(t-4))=1

t/(t-4)=e^1

t/(t-4)=e

t=e(t-4)

t=et-4e

t-et = -4e

t(1-e) = -4e

t = -4e/(1-e)

So the solution is t = -4e/(1-e)